# ZeroCAP: Zero-Shot Multi-Robot Context Aware Pattern Formation via Large Language Models

### v1

@@ Abstract
Incorporating language comprehension into robotic operations unlocks significant advancements in robotics, but also presents distinct challenges, particularly in executing spatially oriented tasks like pattern formation. This paper introduces ZeroCAP, a novel system that integrates large language models with multi-robot systems for zero-shot context aware pattern formation. Grounded in the principles of language-conditioned robotics, ZeroCAP leverages the interpretative power of language models to translate natural language instructions into actionable robotic configurations. This approach combines the synergy of vision-language models, cutting-edge segmentation techniques and shape descriptors, enabling the realization of complex, context-driven pattern formations in the realm of multi robot coordination. Through extensive experiments, we demonstrate the systems proficiency in executing complex context aware pattern formations across a spectrum of tasks, from surrounding and caging objects to infilling regions. This not only validates the system’s capability to interpret and implement intricate context-driven tasks but also underscores its adaptability and effectiveness across varied environments and scenarios.
@@ I Introduction
Pattern formation, a key idea in the organization of multi-robot systems [1], involves arranging robots in a specific way to reach a common objective. This skill is essential for various purposes such as environmental surveillance, precise farming, cooperative transportation, and emergency management. The significance of pattern formation lies in its capacity to utilize the combined actions of robots to carry out tasks that go beyond what individual robots can achieve, thus improving effectiveness, adaptability, and robustness. Central to this endeavor is not just the geometric arrangement of robots, but the evolution towards purposeful pattern formations—where the design and execution of patterns are intrinsically linked to the context and objects within the environment. This nuanced approach marks a significant departure from traditional methodologies, steering the field towards more interactive and object-centric tasks that can be applicable to real world scenarios as shown in Fig 1.
Despite its significance, traditional approaches to pattern formation in robotics have been predominantly constrained by model/rule-based methods and machine learning paradigms. These methodologies, while foundational, have their limitations: model rule-based strategies [1, 2, 3, 4], are often tailored to specific scenarios/patterns and lack the flexibility to adapt to new or unmodeled patterns. Reinforcement learning (RL) approaches, on the other hand, require training time and extensive reward design, making them impractical for scenarios requiring immediate deployment. Moreover, machine learning (ML) based methods [5, 6, 7] demand large datasets of annotated or labeled data for training, a requirement that poses substantial challenges in dynamic or novel environments. These limitations across all the methodologies are particularly pronounced in scenarios where pattern formations must be conditioned with respect to specific objects of interest, underscoring the need for a more adaptable and intuitive methodology.
@@
Integrating language comprehension into robotic systems heralds considerable progress [8, 9], yet it poses unique challenges, especially in spatially-focused tasks such as pattern formation. While language models have shown a remarkable ability as general pattern machines [10] and offer a robust framework for parsing and generating human-like text, Vision Language Models (VLMs) falter in accurately localizing spatial tasks, accentuating the need for techniques that adeptly marry abstract instruction understanding with their tangible implementation.
Our contribution to this evolving field addresses the challenge previously identified by [3], focusing on the dynamic assignment of a variable number of robotic agents to specific target positions for intricate pattern formations. By combining vision-based processing with the sophisticated interpretative capabilities of Large Language Models (LLMs), we introduce a Zero-Shot multi-robot Context Aware Pattern (ZeroCAP) formation system that facilitates zero-shot, language, and context-conditioned pattern formation. It processes both images and natural language instructions to identify objects of interest and generate precise coordinates for robot deployment, enabling immediate and purposeful pattern formation directly influenced by user instructions.
We distinguish our approach by categorizing pattern formations into three purposeful types: 1) General pattern formation, where robots arrange themselves around the object of interest based on user specification; 2) Infill pattern formation, where robots occupy the interior space of a region based on user specification; and 3) Caging pattern formation, creating a boundary around the object. This categorization increases the flexibility of deployment and signifies a shift from standard geometric patterns to contextually relevant formations tailored to the specific user context. Introducing purposeful pattern formation, our system adapts to environmental cues through natural language instructions, demonstrating language and context-driven pattern formation’s viability.
The following sections will explore the technical challenges addressed, methodologies applied, and the significant impact our work has on advancing robotics towards intuitive, flexible, and context-aware multi-robot systems.
@@ II Related Works
Pattern Formation. Multi-robot pattern formation research has explored centralized control [11], [12] and distributed approaches [1], [13]. Centralized systems face scalability and adaptability challenges, while distributed methods, including consensus-based [14], leader-follower [15], and bio-inspired [16], improve scalability but lack robustness due to pattern-specific, rule-based strategies. Advancements to the feild include potential field-based methods[17], [4] and artistic formations [3], plus novel algorithms such as event-triggered mechanisms [2]. However, model/rule-based strategies typically are designed for specific scenarios, lacking the adaptability required for new or unseen patterns or even handling user specific instructions. Conversely, learning-based strategies like Q-learning [7], [6]while offering more generalizability are often time intensive with their training and are limited to the fixed patterns that the system was trained for. Despite the above progress over the years, significant challenges persist in creating flexible solutions that quickly adapt to contextual user preferences and situations without heavy pre-configuration, training, or extensive information exchange. These issues highlight a broader need in the feild for moving from traditional geometric patterns to contextually relevant formations tailored to context specific user tasks and environments.
Our approach diverges from these traditional paradigms by adopting zero-shot learning using a language model for robotics, sidestepping the limitations of both rule-based and machine learning methodologies. By leveraging natural language instructions, our system achieves a flexible, intuitive means of directing multi-robot pattern formation, capable of adaptation to novel scenarios without the need for extensive dataset training or predefined rules. This represents an advancement over existing methods centralized or decentralized.
LLM for Robotics. LLMs are pivotal in robotics for task planning through zero-shot learning [18, 19, 10], excelling in diverse plan generation [20, 9, 8] and performance refinement [21]. In the realm of multi-agent systems, LLMs have been instrumental in fostering agent cooperation, equipping individual agents with the capabilities to improve interaction dynamics and collaborative efficacy [22, 23, 24]. Despite their growing utility in robotics, none of these noteable methodologies cover pattern formation among multi-robot systems. Vision Language Models, while useful in many applications involving robot navigation[25] and high-level action control[26, 27], are limited by their spatial reasoning capabilities [28], impacting their effectiveness in robotics applications that require precise spatial coordination. Our research leverages VLMs and LLMs for multi-robot pattern formation, addressing the spatial task limitations of VLMs and pioneering a linguistically-driven approach to complex robotic coordination aided by representing shapes using a node graph.
@@ III Problem Description
Our research delves into the nuanced task of deploying robotic agents to form patterns around or inside a designated object of interest $O$, located within an environment, $e$. This environment is depicted through an image ($I_{E}$), paired with a natural language instruction ($Ins$) that outlines the desired pattern formation task. Operating under a centralized control system, robots, simplified to points of mass, are coordinated to ensure precise pattern formation in the context of $O$ and $Ins$.
@@
Let $R=\{r_{1},r_{2},\ldots,r_{N}\}$ represent the list of robotic agents available, where $N$ is the total number of robots. Given $Ins$, and $I_{E}$, the system, through a series of steps that include object detection, segmentation, and shape determination resulting in $G$, leverages an LLM to interpret and reason natural language instructions in light of $G$. The output of the system is a list of $(x,y)$ coordinates, $C_{xy}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$, which assigns each robot $r_{i}$ in $R$ to a specific coordinate $(x_{i},y_{i})$ for deployment. This assignment facilitates the formation of the specified pattern around/inside $O$, in accordance with $Ins$.
Formally, this process can be encapsulated as a function $F$, where $F(I_{E},Ins)\rightarrow C_{xy}$, mapping the inputs to the deployment coordinates for each robot.
Figure 2: An overview of the ZeroCAP system. It traces the workflow from the initial natural language instruction and input image of the environment through to the final deployment of robots, illustrating the sequence of processing stages—including context identification using Vision Language Model (VLM), object segmentation, shape description, and Large Language Model (LLM) coordination for precise robot placement in the environment. Three key stages are highlighted and explained in Section IV.
@@ IV ZeroCAP Formation System
This section delineates the comprehensive methodology of our ZeroCAP formation system, which introduces a novel approach for zero-shot multi-robot context aware pattern formation. Leveraging advanced VLMs, the methodology integrates linguistic instructions with visual perception, enabling precise spatial reasoning and object identification in multi-robot systems. Our approach harnesses the synergy of cutting-edge segmentation techniques, and utilizes standard computer vision algorithms for obtaining shape description. The decision-making process is empowered by a specialized LLM, which synthesizes the spatial description of object and linguistic context to generate target deployment coordinates for the robotic agents. The multi-robot control is centrally orchestrated, ensuring efficient navigation. The following subsections provide a detailed exposition of each stage, from initial input processing to the final deployment of robots.The overall structure of the proposed ZeroCAP framework is shown in Fig. 2.
@@ IV-A Context Identification using VLM
The initiation of our methodology engages a VLM to interpret an environmental image $I_{E}$ and a natural language instruction $Ins$. The architecture of this model excels at correlating visual stimuli with corresponding linguistic descriptors, emulating human-like proficiency in image-language tasks. Provided with $I_{E}$ and $Ins$, the VLM processes these inputs in tandem, producing two key outputs: the Object of Interest $O$ and the Pattern Formation Instruction $PI$.
The Object of Interest $O$ is identified utilizing the VLM’s zero-shot detection capabilities, circumventing the need for prior object-specific training and ensuring versatility within varying environments. Conventional object detection methods may localize objects but lack the sophisticated reasoning to discern context as VLMs do. Given an instruction akin to that of Fig. 3, standard natural language-aided object detectors [5] would falter in reasoning with which object is odd. In contrast, the advanced reasoning capabilities of the VLM, contextualized by the instruction, render it superior for identifying objects of interest within complex scenarios. This nuanced understanding of context allows the VLM not only to detect but also to reason about objects in a manner aligned with the given linguistic directives. Concurrently, the Pattern Formation Instruction $PI$ is extrapolated from the natural language directive enriched by contextual nuances in $I_{E}$. Here, the VLM generates $PI$, inspired by chain-of-thought reasoning [30], to discern spatial and context-sensitive task directives. In Algorithm 1, Line 1 corresponds to estimating the context using the VLM.
One notable strength of our system is our approach to mitigating the common issue of precise spatial localization, a challenge in standard VLMs. The inherent limitations of VLMs in tasks that require exact positional outputs make it impractical to use these models to directly provide the deployment coordinates $C_{xy}$. Instead, the VLM is optimized to identify $O$ with necessary precision and to extract a context-rich $PI$. These are indispensable for subsequent segmentation and precise pattern formation stages conducted by more specialized subsystems. By honing in on general features and attributes discernible in the environment and the language input, the VLM identifies $O$ with the necessary precision for subsequent segmentation and pattern formation stages.
The VLM’s output, consisting of $O$ and $PI$, thus propels the ZeroCAP system into the segmentation phase, where $O$ is visually distinguished within $I_{E}$, and $PI$ informs the strategic pattern formation instruction to be implemented by the robotic entities.
@@ IV-B Object Segmentation and Shape Description
Following the identification of $O$ and the acquisition of the $PI$ from the VLM, the system transitions into the crucial stages of segmentation and shape description. These stages leverage sophisticated language conditioned segmentation techniques [5, 31], to perform image segmentation, followed by shape description stage using standard computer vision methods.
@@ IV-B1 Object Segmentation
The segmentation process begins with language-prompt-based object segmentation, which employs LangSAM, an open-source language-conditioned segmentation model built upon segment-anything approach and the GroundingDINO detection model [31, 32]. Provided with the Object of Interest $O$ and the Input Image $I_{E}$, these models generate a binary segmentation map in which the object is highlighted in white against a black background.
Upon completion, the segmentation yields a Masked Image $I_{M}$ and an accuracy metric $\alpha$, compared against a predefined threshold $t$. If $\alpha$ meets or exceeds this threshold, the process proceeds, confirming the segmentation’s quality for subsequent tasks. This segmentation approach ensures the precise extraction of $O$ from $I_{E}$, encapsulated within $I_{M}$ for the next processing stage. Lines 2 to 4 of the Algorithm 1 expound the segmentation process.
@@ IV-B2 Shape Description
With a high quality $I_{M}$ at hand, the ZeroCAP system advances to the shape description phase. Here, standard computer vision techniques preprocess $I_{M}$, including filtering and enhancement operations to sharpen feature delineation. Through OpenCV’s edge detection algorithms and contour detection methods, $I_{M}$ is analyzed to determine $O$ geometric form. This analysis yields the object’s edges $E$ and vertices $V$, forming a connected graph $G=(V,E)$ that mathematically represents the object’s structure. Algorithm 1 delineates this process in Lines 10 – 13, with a method call in Line 5.
Integrating these segmentation techniques with precise shape description, the ZeroCAP system ensures accurate recognition, segmentation, and modeling of the Object of Interest. This forms a solid foundation for the ensuing methodology phase: LLM-based Pattern Former.
@@
9:for each coordinate $(x_{i},y_{i})$ in $C_{xy}$ do
10:   $r_{i}\leftarrow R[i]$
11:   DeployRobots($r_{i},(x_{i},y_{i})$)
12:procedure ShapeDescriptor($I_{M}$)
13:   Preprocess the masked image $I_{M}$
14:   Extract edges $E$ and vertices $V$
15:   return $(V,E)$ from $I_{M}$
16:procedure DeployRobots($r,(x,y)$)
17:   Command robot $r$ to move to coordinates $(x,y)$
18:   Verify that robot $r$ has reached $(x,y)$

### v2

@@ Abstract
Incorporating language comprehension into robotic operations unlocks significant advancements in robotics, but also presents distinct challenges, particularly in executing spatially oriented tasks like pattern formation. This paper introduces ZeroCAP, a novel system that integrates large language models with multi-robot systems for zero-shot context aware pattern formation. Grounded in the principles of language-conditioned robotics, ZeroCAP leverages the interpretative power of language models to translate natural language instructions into actionable robotic configurations. This approach combines the synergy of vision-language models, cutting-edge segmentation techniques and shape descriptors, enabling the realization of complex, context-driven pattern formations in the realm of multi robot coordination. Through extensive experiments, we demonstrate the systems proficiency in executing complex context aware pattern formations across a spectrum of tasks, from surrounding and caging objects to infilling regions. This not only validates the system’s capability to interpret and implement intricate context-driven tasks but also underscores its adaptability and effectiveness across varied environments and scenarios. The experimental videos and additional information about this work can be found at https://sites.google.com/view/zerocap/home.
@@ I Introduction
Pattern formation, a key idea in the organization of multi-robot systems (MRS) [1], involves arranging robots in a specific way to reach a common objective. This ability is critical for applications such as environmental surveillance, precision farming, cooperative transportation, and emergency management. Its value lies in coordinating robots to perform tasks beyond the capabilities of individuals, enhancing effectiveness, adaptability, and robustness. In this paper, we go beyond traditional geometric arrangements, focusing on purposeful pattern formations-where patterns are closely tied to user provided contexts and environments. This approach marks an evolution in MRS, steering the field towards more interactive and object-centric tasks that can be applicable to real world scenarios as shown in Fig 1.
Despite its significance, traditional approaches to pattern formation in robotics have been predominantly constrained by model/rule-based methods and machine learning paradigms. These methodologies, while foundational, have their limitations: model/rule-based strategies [1, 2, 3, 4] are often scenario-specific and lack adaptability to new patterns, while reinforcement learning (RL) requires significant training and reward design, making it impractical for immediate deployment. Moreover, machine learning (ML) methods [5, 6, 7] rely on large, annotated datasets, which are challenging in dynamic environments. These limitations are especially evident when pattern formation must be conditioned on user-provided natural language inputs referring to specific objects, emphasizing the need for a more flexible and intuitive approach.
@@
Integrating language comprehension into robotic systems offers considerable progress [8, 9], but it also presents challenges in spatial tasks like pattern formation. While language models have shown a remarkable ability as general pattern machines [10] and offer a robust framework for parsing and generating human-like text, Vision Language Models (VLM) face challenges in precise spatial localization tasks due to limitations in spatial reasoning and dataset training [11, 12]. This underscores the need for approaches that combine abstract instruction interpretation with accurate spatial execution by robots.
Our contribution to this evolving field addresses the challenge previously identified by [3], focusing on the dynamic assignment of a variable number of robotic agents to specific target positions for intricate pattern formations. By combining vision-based processing with the sophisticated interpretative capabilities of Large Language Models (LLMs), we introduce a Zero-Shot multi-robot Context Aware Pattern (ZeroCAP) formation system that facilitates zero-shot, language, and context-conditioned pattern formation without any prior training or data. It processes both images and natural language instructions to identify objects of interest and generate precise coordinates for robot deployment, enabling immediate and purposeful pattern formation directly influenced by user instructions.
The main contributions of this paper are as follows:
  • ZeroCAP, a novel zero-shot multi-robot pattern formation framework that integrates LLMs with vision-based processing to translate user natural language inputs into context-aware, purposeful pattern formations, moving beyond traditional geometric arrangements.
  • Implementation and evaluation of the framework in both simulation and real-world environments, validating the system’s ability to perform complex pattern formations with varied objects and natural language inputs.
The following sections will explore the challenges tackled, methodologies employed, and the broader implications of our work for advancing intuitive, flexible, and context-aware MRS.
@@ II Related Works
Pattern Formation. Multi-robot pattern formation has been explored through both centralized [13], [14] and distributed methods [1], [15]. Centralized systems, while efficient in task coordination, can suffer from single points of failure and may be less adaptable in highly dynamic environments. Distributed approaches, such as consensus-based [16], leader-follower [17], and bio-inspired methods [18], improve adaptability and eliminate single points of failure but often rely on rule-based strategies that lack flexibility when applied to diverse or complex tasks. Advancements like potential field methods [19], [4], and event-triggered mechanisms [2] have been proposed, yet rule-based models still lack adaptability to novel or unseen patterns, while approaches like Mora et al. [3], emphasize visually appealing, artistic robot movements. Learning-based strategies like Q-learning [7], [6], while more generalizable, require extensive training and are limited to fixed patterns. Despite progress, challenges remain in developing flexible solutions that quickly adapt to user-specific contexts without pre-configuration or heavy training, moving beyond traditional geometric patterns to more contextually relevant formations.
Our approach bypasses these limitations by employing zero-shot learning with a language model, enabling flexible, intuitive multi-robot pattern formation. Leveraging natural language instructions, our system adapts to novel static scenarios without extensive training or predefined rules, marking a significant improvement over existing methods.
LLM for Robotics. LLMs are pivotal in robotics for task planning through zero-shot learning [20, 21, 10], excelling in diverse plan generation [22, 9, 8] and performance refinement [23]. In the realm of multi-agent systems, LLMs have been instrumental in fostering agent cooperation, equipping individual agents with the capabilities to improve interaction dynamics and collaborative efficacy [24, 25, 26]. Despite their growing utility in robotics, none of these noteable methodologies cover pattern formation among MRS.
VLM, while useful in many applications involving robot navigation[27] and high-level action control[28, 29], are limited by their spatial reasoning capabilities [12], impacting their effectiveness in robotics applications that require precise spatial coordination. Our research leverages VLMs and LLMs for multi-robot pattern formation, addressing the spatial task limitations of VLMs and pioneering a natural language-driven approach to complex robotic coordination aided by representing shapes using a node graph.
@@ III Problem Description
This research focuses on the task of deploying robotic agents to form patterns around or inside a designated object of interest ($O$) within an environment ($e$), depicted through an image ($I_{E}$) and guided by a natural language instruction ($Ins$). We operate under a centralized control system where robots are simplified as point masses and coordinated to achieve precise pattern formation based on $O$ and $Ins$. This centralized approach is adopted to streamline task coordination by eliminating the need for each robot to independently process language and spatial data. It reduces computational overhead and avoids the significant costs of equipping each robot with its own VLM/LLM.
@@
Let $R=\{r_{1},r_{2},\ldots,r_{N}\}$ represent the list of robotic agents available, where $N$ is the total number of robots. Given $Ins$, and $I_{E}$, the system, through a series of steps that include object detection, segmentation, and shape determination resulting in $G$, leverages an LLM to interpret and reason natural language instructions in light of $G$. The output of the system is a list of $(x,y)$ coordinates, $C_{xy}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$, which assigns each robot $r_{i}$ in $R$ to a specific coordinate $(x_{i},y_{i})$ for deployment. This assignment facilitates the formation of the specified pattern around/inside $O$, in accordance with $Ins$. While the current formulation operates in 2D, based on the 2D input image $I_{E}$, it can be generalized to n-dimensions.
Formally, this process can be encapsulated as a function $F$, where $F(I_{E},Ins)\rightarrow C_{xy}$, mapping the inputs to the deployment coordinates for each robot. This function represents the ZeroCAP system, which is detailed further in Sec. IV.
Figure 2: An overview of the ZeroCAP system. It traces the workflow from the initial natural language instruction and input image of the environment, to the final deployment of robots, illustrating the sequence of processing stages—including context identification using Vision Language Model (VLM), object segmentation, shape description, and Large Language Model (LLM) coordination for precise robot placement in the environment. Three key stages are highlighted and explained in Section IV.
@@ IV ZeroCAP Formation System
This section outlines the methodology behind the ZeroCAP formation system. The system facilitates zero-shot, context-aware pattern formation for MRS by integrating LLMs with visual perception and analysis. The system, represented by the function $F$, takes two key inputs: $I_{E}$ and $Ins$. The VLM processes these inputs to identify $O$ and generate a pattern formation instruction ($PI$). The system then segments the object, extracts shape descriptors, and the LLM uses this spatial descriptors along with the instruction to generate precise deployment coordinates for the robots. Finally, under centralized control, the robots execute the task based on the generated coordinates. The overall structure of the ZeroCAP framework is depicted in Fig. 2. Algorithm 1 outlines the execution pipeline of the system, from processing the inputs to deploying robots to the designated coordinates.
@@ IV-A Context Identification using VLM
This stage, represented as box ‘Sec IV. A’ in Fig. 2, begins with a VLM interpreting the environmental image $I_{E}$ and $Ins$. The VLM is proficient at correlating visual and linguistic inputs, emulating human-like proficiency in image-language tasks. This stage produces two key outputs: $O$ and $PI$.
$O$ is detected using the VLM’s zero-shot capabilities, eliminating the need for prior object-specific training and allowing versatility in diverse environments. Unlike conventional object detection, which may localize objects but lacks contextual reasoning, the VLM can reason about the context and identify complex instructions [12]. Taking the case of Fig. 3, standard natural language-aided object detectors [5] struggle to reason about which object is odd, whereas the VLM’s advanced reasoning, guided by the user instruction, makes it more effective at identifying objects in complex scenarios. This advanced reasoning enables the VLM to align object identification with the given language directives.
Simultaneously, the $PI$ is derived from the natural language directive, enriched by the context in $I_{E}$. Here, the VLM generates $PI$, inspired by chain-of-thought reasoning [31], to discern spatial and context-sensitive task directives. Algorithm 1, Line 1 corresponds to estimating the context using the VLM.
One key strength of our system is addressing the challenge of precise spatial localization, a known limitation of standard VLMs. VLMs are not well-suited to directly generate $C_{xy}$, so instead, they focus on accurately identifying the $O$ and extracting a context-rich $PI$. These outputs are crucial for subsequent segmentation and pattern formation, handled by more specialized subsystems below. By focusing on general features and context from the environment and language input, the VLM ensures accurate identification of $O$ for further processing.
The VLM’s output, consisting of $O$ and $PI$, thus propels the ZeroCAP system into the segmentation phase, where $O$ is visually distinguished within $I_{E}$, and $PI$ informs the strategic pattern formation instruction to be implemented.
@@ IV-B Object Segmentation and Shape Description
Following the identification of $O$ and acquiring the $PI$ from the VLM, the system moves to segmentation and shape description represented as box ‘Sec IV. B’ in Fig. 2. These stages utilize language-conditioned segmentation [5, 32], and standard computer vision methods for shape analysis.
@@ IV-B1 Object Segmentation
The segmentation process employs LangSAM, an open-source language-conditioned model based on the segment-anything approach and GroundingDINO [32, 33]. Given $O$ and $I_{E}$, the model generates a binary segmentation map, highlighting the object in white against a black background. The segmentation outputs a Masked Image $I_{M}$ and an accuracy metric $\alpha$, which is a scalar value compared against a threshold $t$, a predefined scalar constant. If $\alpha$ exceeds $t$, the process continues, ensuring the segmentation quality is sufficient. This approach precisely extracts $O$ from $I_{E}$, encapsulated in $I_{M}$ for further processing (Algorithm 1, Lines 2-4).
@@ IV-B2 Shape Description
With a quality $I_{M}$, the system proceeds to shape description. Using standard computer vision techniques, $I_{M}$ is preprocessed for edge and contour detection through OpenCV, resulting in the identification of edges $E$ and vertices $V$. These form a connected graph $G=(V,E)$, which mathematically represents $O$. The process is shown in Algorithm 1, Lines 9-12, initiated by Line 5.
ZeroCAP thus integrates segmentation and shape description to ensure accurate object recognition and modeling, forming the basis for the LLM-based Pattern Former.
@@
9:for each robot $r_{i}$ in $R$ and coordinate $(x_{i},y_{i})$ in $C_{xy}$ do
10:   DeployRobots($r_{i},(x_{i},y_{i})$)
11:procedure ShapeDescriptor($I_{M}$)
12:   Preprocess the masked image $I_{M}$
13:   Extract edges $E$ and vertices $V$
14:   return $(V,E)$ from $I_{M}$
15:procedure DeployRobots($r,(x,y)$)
16:   Command robot $r$ to move to coordinates $(x,y)$
17:   Verify that robot $r$ has reached $(x,y)$